A car moving along a straight line (without returning) travels distance equal to the numerical value of time elapsed in every second. If car starts at t = 0 with constant acceleration, the magnitude of initial velocity of the car is 
 
1) 0.2 m/s
2) 0.5 m/s
3) 0.75 m/s
4) 1.0 m/s

Grade:12th pass

1 Answers

Susmita
5 years ago
Let it start with velocity u and acceleration a.
For 1st second:
v=u+at=u+a
s=(v2-u2)/2a=u+a/2=1
For the 2nd second:
v'=(u+a)t+at=u+2a
s'=u+3a/2=u+a/2+a=1+a=2(s'=2 as 2s has elasped).
So 1+a=2
Or,a=1.
Putting it in the equation of s,
u+a/2=1
Or,u=1/2=0.5
If it helps please approve.